OpenStudy (anonymous):

Express the function as the sum of a power series by first using partial fractions and find the interval of convergence. See comments for the function

OpenStudy (anonymous):

\[f(x)=\frac{ x+9 }{ 2x^2-15x-8 }\]

OpenStudy (anonymous):

\[ \frac{-1}{2x+1} + \frac{1}{x-8} =\frac{\frac{-1}{8}}{1-\frac{x}{8}}+\frac{1}{-2x-1} \]